THE ROTATIONAL SPECTRA OF S_{3} AND S_{4}

\underline{C. A. GOTTLIEB}, M. C. McCARTHY, S. THORWIRTH, and P. THADDEUS, Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138, U.S.A. and Division of Engineering and Applied Sciences, Harvard University, 29 Oxford Street, Cambridge, MA 02138, U.S.A..

Following the initial detection of the pure sulfur clusters S_{3} and S_{4} at centimeter wavelengths by Fourier transform microwave spectroscopy^{a}, S_{3} has subsequently been observed at millimeter wavelengths. Thirteen spectroscopic constants reproduce over 60 transitions of S_{3} between 9 and 460 GHz (with J ≤ 87 and K_{a} ≤ 7) to within the measurement uncertainties. From these the frequencies of the astrophysically relevant lines of S_{3} can be calculated to about 1 part in 10^{7} up to 500 GHz, allowing deep searches in the atmosphere of Io, dense molecular cores, and circumstellar shells of late-type stars with existing millimeter- and submillimeter-wave telescopes.

aM. C. McCarthy, S. Thorwirth, C. A. Gottlieb, and P. Thaddeus, J. Am. Chem. Soc., in press.